OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Seven divided by x squared. + 10

OpenStudy (anonymous):

\[\frac{ 7 }{ x ^{2}+10 }\] So this is how it looks like?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

well f(g(x) is a composite function. This meaning that there is a function INSIDE of anpther function. So in this case we looking for a function g INSIDE of f. Does this make sense so far?

OpenStudy (anonymous):

So what we are trying to do in see a function inside of a function.. With this we can see that f(x)=the outside function and g(x is the inside function. so f(x)= 7/x g(x)=x^2=10. Why? because here g is inside of f... If we replace all the x's in the f(x) function with the g(x) function we will get our y function of f(g(x)).

OpenStudy (anonymous):

I hope that makes sense

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Youre very welcome!